Please show all workings and steps for question 3:

2 3. Determine the area of the region bounded by y = 12 – 2x2 and y = x? – 9.
Answer #1


(given Y= 12-222 yox?_g point of inteosechori 12-272, 12-9 21 and = 3x2. T = Iva 27 Area undeo cureve = 12-282) - (x² - g) da

Please do like